OpenStudy (anonymous):

Tom borrowed $35,000 to remodel his house. At the end of the 5-year loan, he had repaid a total of $46,375. At what simple interest rate did he borrow the money?

OpenStudy (anonymous):

total amount paid = p * i * n p = principal (original amount borrowed or loaned) i = interest rate for one period n = number of periods 46,375 = 35,000*5*x 46,375 = 175000x 46,376/175000 = x x = .265 times the .265 by 100 = 26.5%

OpenStudy (anonymous):

Okay.Thanks.:)